### v1

@@ 1 Introduction
Most recently, Evidential Physics-Informed Neural Networks (EPINN) was proposed in [22], where principles of Evidential Deep Learning (EDL) [^1] Evidential Deep Learning [2, 21] is a framework where a higher-order set of priors are placed over likelihood functions for the outputs, and model training is then used to infer uncertainty estimates from closed-form functions of the learned prior hyperparameters. was blended with those of PINN to yield uncertainty estimates for the target outputs assumed to follow a set of PDEs. EPINN [22] was conjectured to be able to generate uncertainty estimates also for the unknown PDE parameters (apart from the target output) by averaging the PDE residual term over a hypothetical distribution. A limitation of the work of [22] is that it only applies to a subclass of PDEs where its parameters can be separated out as linear terms in a certain form (as expressed in equation 3 of [22]). It was also not clear how one should specify the hypothetical distribution for the learnable PDE parameters.
@@ 2.1 A posterior density from merging principles of Deep Evidential Regression and PINN
where $N_{D}$ is the number of independently sampled points within the domain of the PDE and $\mathcal{R}\left(\partial y,y,x,\vec{\Omega}\right)=0$ refers to the partial differential equation, with $\sigma^{2}_{R}$ being the variance parameter of a density function for the residual. [^2] One can extend this to a system of PDEs by defining a separate residual for each differential equation. The negative logarithm of (2) is the PDE residual loss, with the parameter $\sigma^{2}_{R}$ quantifying the weight of this loss term relative to the data loss term $\sim-\log P(\mathcal{D}|\mathcal{M}(\vec{w}))$. In standard PINN, this is a free parameter and, to our knowledge, there is no principled approach towards determining its choice. Here we lift $\sigma^{2}_{R}$ to be a learnable parameter with the relative weight of the PDE residual evolving as the model shifts towards a minimum in the loss landscape. We regularize the dynamical evolution of $\sigma^{2}_{R}$ through a prior density function $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$ which we pick to be the inverse-gamma distribution. Since the residual is a function of the unknown parameters $\vec{\Omega}$, the form of (1) suggests that we can interpret it as the (unnormalized) posterior density for $\vec{\Omega}$, with $\pi(\vec{\Omega})$ being the prior distribution and the product $P(\mathcal{D}|\mathcal{M}(\vec{w}))\,P(\mathcal{M}(\vec{w})|\vec{\Omega})$ being the likelihood function. The model’s weights are then latent variables with model training equivalent to a maximum a posteriori (MAP) estimation. This interpretation enables us to compute the uncertainty of $\vec{\Omega}$ as being defined with respect to the posterior density function
@@
where $\gamma$ is the mean output and $\alpha,\beta,\nu$ are learnable functions of input related to the overall predictive uncertainty $\sigma^{2}_{p}$ as follows.
@@
where $Y$ denotes each output variable in the dataset $\mathcal{D}$, and the expectation and variance are taken with respect to the t-distribution density function in (6). This distribution is the marginal distribution obtained after integrating the mean and variance of a Gaussian likelihood function $\mathcal{N}(\mu,\sigma^{2})$ with respect to a normal-inverse-gamma prior distribution, where $\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\nu)$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$. We take the dependent variable in the PDE residual to be the mean output, i.e. $y\rightarrow\gamma(x)$ in eqn. (2). This implies that the MAP estimation involves the adherence of the mean target variables to the PDEs. Upon completion of model training, we can place confidence intervals on model’s predictions using the learned function $\sigma^{2}_{p}$. Since we infer $\vec{\Omega}$ at the end of model training via eqn. (3), our framework thus appears as a maximum a posteriori estimation of $\vec{\Omega}$, or more precisely a MLE estimation that is regularized by the prior $\pi(\vec{\Omega})$.
@@ 2.2 On uncertainty of $\vec{\Omega}$ and its prior distribution
In our framework, we alluded to a posterior density function $f_{p}\left(\vec{\Omega}|\mathcal{D},\mathcal{M}(\vec{w})\right)$ in eqn. (3) of which negative logarithm is the model’s loss function. Considering the nature of eqn.(3), the uncertainty in $\vec{\Omega}$ is fundamentally defined as the degree of deviation of the model from the PDE description, as measured by the residual. In the following, we will invoke this principle to derive a form for the prior $\pi(\vec{\Omega})$. Let $\mathcal{D}_{\vec{\Omega}}$ denote the finite, discretized domain for the unknown parameters $\vec{\Omega}$. At each point of $\mathcal{D}_{\vec{\Omega}}$, we can evaluate the mean squared deviation between the numerical solution to the PDE characterized by $\vec{\Omega}$ ($L_{p}(x;\vec{\Omega})$) and the data-fitted model $\mathcal{M}_{0}(x;\vec{w}_{0})$.
@@
where $N$ is the normalization constant and all integrals are implemented as numerical Riemann sums over the discretized domain $\mathcal{D}_{\vec{\Omega}}$. We would like the prior distribution of $\Omega$ to be characterized by the same mode and dispersion scales as the highest density region (see e.g. [14]) of $f$. At some confidence level, say 68%, this region is generally a complex subset of the domain $\mathcal{D}_{\Omega}$. Since our choice of prior distribution affects model training dynamics in the second phase, we adopt a simple Gaussian surrogate distribution for this region, with the means being the modes and the standard deviations being those of each marginal distribution.
@@ 2.3 Determination of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$
we pick its hyperparameters $(\alpha_{r},\beta_{r})$ such that it is consistent with other aspects of our formalism. The mode and mean values are $\frac{\beta_{r}}{\alpha_{r}+1}$ and $\frac{\beta_{r}}{\alpha_{r}-1}$ respectively. Here we restrict ourselves to the case where $\alpha_{r}>1$ so that the mean is well-defined. We pick the initial value of $\sigma^{2}_{R}$ ($\equiv\sigma^{2}_{ini}$) to be the mean. As $\sigma^{2}_{R}$ decreases during model training, it approaches the mode of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$ at which the derivative with respect to $\sigma^{2}_{R}$ vanishes.
@@ 3 Validation case studies: 1D Poisson and 2D Fisher-KPP equations
We validated the E-PINN framework using two representative differential equations : (i)the 1D Poisson equation with Gaussian source (ii)the Fisher-KPP equation. These prototype equations and their variants have found wide applicability across scientific disciplines (see e.g. [10, 13, 6, 19].) For the main performance metric, we adopt the empirical coverage probability (ECP) to measure how well-calibrated the uncertainties are, relative to the nominal ones (NCP). In our case studies, we computed plots of ECP vs NCP and calculated the mean of the absolute discrepancy between these two quantities as a measure of the mean calibration error (MCE) (see e.g. [15, 12]). For assessing if the uncertainty in the inferred parameters $\vec{\Omega}$ is consistent with the model’s deviation from the PDE numerical solutions, we performed a Monte Carlo goodness-of-fit test (see e.g. [24]). This involves estimating an empirical p-value defined as the proportion of model deviations smaller than the simulated solutions’ deviations (from the mean of simulated set).
@@
where $x_{0},\sigma^{2}_{f}$ are parameters to be inferred from the model. The synthetic dataset was based on an exact solution $u_{e}(x)$ defined on the finite segment $x\in[0,1]$, with $x_{0}=1/3,\sigma^{2}_{f}=0.02$. Dirichlet boundary conditions $u=0$ are imposed at $x=0,1$. A uniformly distributed random noise of the form $\lambda_{N}\,\mathcal{U}(-0.5,0.5)$ was added to the exact solution within $(0,1)$. We chose $\lambda_{N}$ to be 0.1 of the maximum value of $u_{e}$. The size of the training and testing data was taken to be 240 and 150 (randomly chosen) points respectively covering the entire finite domain. [^3] About half of the training dataset was used as a validation set for choosing the optimal learning rate in the first phase of model training, after which the whole training dataset was used.
For this problem, our base model was a fully-connected two-layer perceptron model with 16 neurons per layer. In the first training phase where the model was fitted purely to data, convergence was attained upon $2\times 10^{5}$ epochs at a learning rate of $10^{-4}$, with the Adam optimizer implemented via PyTorch. Using this initial model, we determined the prior distribution $\pi(\vec{\Omega})$ following our earlier discussion in Section 2.2. After marginalizing each parameter to obtain estimates of the dispersion scales associated with this HDR, we constructed the prior distribution $\pi(\vec{\Omega}$ following eqn. (10), (11) (see the yellow region in Fig. LABEL:fig:prior). Once the prior distribution $\pi(\vec{\Omega})$ was determined, we used it to determine the prior for $\sigma^{2}_{R}$ using eqn.15-18. The finite parameter domains were set to be $x_{0}\in[0,1],\sigma^{2}_{f}\in[0.01,0.06]$ discretized such that the minimal spacing is 1/50 of each interval. We then proceeded to the second phase of model training with a learning rate of $5\times 10^{-4}$. Upon completion of training, our model inferred the equation’s parameters to be
@@
where the upper and lower figures refer to the 68% confidence intervals. We note that they contain the actual values of these parameters ($x_{0}=0.333,\sigma^{2}_{f}=0.02$). The p-value for the Monte Carlo goodness-of-fit test was 0.89, showing lack of evidence against the hypothesis that the model curve arises from the inferred posterior distribution characterized by (19). The model’s uncertainty distribution appeared to be well-calibrated, with the ECP close to the NCP for all confidence intervals. The mean calibration error was 0.02. Figure 1 shows the model-predicted curve and the ECP vs NCP plot revealing a well-calibrated uncertainty distribution.
@@
We set r = 1.6, D = 6.2 for generating the training dataset, with these values intended to be recovered from the model. Like in the Poisson equation case study, we added a uniformly distributed random noise $\lambda_{N}\mathcal{U}(-0.5,0.5)$ to the exact solution (21) defined on the 2D domain $x\in(-20,20),t\in(0,10)$ (see Fig. 2). The size of the training and testing datasets was set to be $5\times 10^{3}$ and $4\times 10^{3}$ randomly distributed points across the domain. Like in the previous case study, we used a fully connected two-layer perceptron model with 16 neurons per layer. The training dynamics was similar to the previous case, with the model trained for $2\times 10^{5}$ epochs using only the data loss term and the prior distribution constructed from the HDR of the empirical distribution as defined in eqn. (9). The second phase of model training was implemented at $3\times 10^{-4}$ learning rate and convergence reached after $3\times 10^{5}$ epochs. At the end of training, our model inferred the equation’s parameters to be
@@
We note that the 68% confidence intervals in (22) cover the actual values $(r=1.6,D=6.2)$. The p-value for the goodness-of-fit test was 0.93, indicating lack of evidence against the posterior distribution of (22). The model’s uncertainty quantification appeared to be well-calibrated similar to the previous Poisson equation case study, with a mean calibration error of 0.024.
@@ 3.3 Comparison against other methods of uncertainty quantification frameworks in PINN
In terms of the degree of uncertainty calibration, we found E-PINN to perform significantly better than Deep Ensemble and B-PINN models for both case studies. While B-PINN tended to yield over-conservative uncertainty estimates, Deep Ensemble models generated excessively over-confident ones. The ECP vs NCP plots showed E-PINN’s uncertainty estimates to be well-calibrated in contrast to the more obvious deviations from the NCP for the other two frameworks. The MCE values for E-PINN were much lower relative to B-PINN and Deep Ensemble. For the 2D Fisher-KPP case study in which heterogeneous noise was inserted into the training data, we found that the E-PINN uncertainty estimates yield a strong correlation with the training data noise (Spearman coefficient $r_{s}\sim 0.8,p<0.01$) which was absent in Deep Ensemble ($\rho_{s}\sim 0.01,p=0.2$) and B-PINN ($\rho_{s}\sim 0.09,p<.001$) as evident in the Fisher-KPP uncertainty plots (Fig. 3).
@@ 4 Application of E-PINN to Bergman minimal model of glucose-insulin regulatory system
where $G$ represents glucose level, $I$ represents insulin concentration and $X$ a more abstract quantity representing glucose uptake activity of insulin-excitable tissue (e.g. liver, adipose, etc.). $G_{b},I_{b}$ are the basal concentrations of glucose and insulin representing baseline values when the effect of the bolus injection of glucose in IVGTT disappears. The parameters $p_{1},p_{2},p_{3}$ represent the glucose decay rate, spontaneous decay rate of remote insulin and the insulin-dependent increase in tissue glucose uptake ability respectively. These are the parameters we leverage E-PINN to infer from measurements of glucose and insulin evolutions in IVGTT samples. In particular, $p_{1}$ is known as an index measuring ‘glucose effectiveness’ (often denoted as $S_{G}$), while $S_{I}\equiv p_{3}/p_{2}$ is known as the insulin sensitivity index that measures the effectiveness at which glucose decreases upon release of insulin. In a IVGTT test, glucose and insulin samples are periodically collected over a period of a hours. Taking the insulin to be an externally known function, the equations (23), (24) thus constitute a coupled set of first-order ODEs in the variables $G$ and $X$ which is often alluded to as the Bergman minimal model [3].
@@
We apply E-PINN to the modeling of (23) and (24), using the datasets gathered and analyzed in [9, 16, 23]. In particular, we focused on modeling two datasets pertaining to the measured glucose and insulin levels corresponding to healthy and type 2 diabetic subjects. E-PINN enables probabilistic modeling of the datasets and directly yields the posterior distribution of the unknown parameters. Unlike the controlled case studies, we do not have empirical data points for the variable $X(t)$ and thus, our base neural network was designed to yield five outputs – four corresponding to $G(t)$ and its uncertainty and another corresponding to $X(t)$ so as to avoid overfitting on $X(t)$. We define the PDE residual as the sum of the residual arising from (23) and (24) separately. Estimating $I_{b},G_{b}$ from the datasets, we are left with only $p_{1},p_{2},p_{3}$ as the unknown parameters to be inferred. Fig. 4 depicts our model’s prediction curves and the learned posterior distributions of two clinically useful indices: the glucose effectiveness index $S_{G}$ and insulin sensitivity index $S_{I}$, the latter appearing to be the more discriminatory one, being about 3 orders of magnitude larger for healthy individuals.
@@
Figure 4: The left diagrams show E-PINN’s prediction curves for glucose evolution in healthy and diabetic subjects. Blue dashed curves are the numerical solutions to Bergman’s ODE with parameters $p_{1},p_{2},p_{3}$ being the posterior modes. Deviations between blue and red curves in the middle regions suggest that PDE description may be relatively less valid during those durations. The right diagram superposes the learned posterior distributions of the indices $S_{G},S_{I}$, revealing the latter as a more discriminatory index for identifying diabetic individuals.

### v2

@@ 1 Introduction
Most recently, Evidential Physics-Informed Neural Networks (E-PINN) was proposed in [22], where principles of Evidential Deep Learning (EDL) [^1] Evidential Deep Learning [2, 21] is a framework where a higher-order set of priors are placed over likelihood functions for the outputs, and model training is then used to infer uncertainty estimates from closed-form functions of the learned prior hyperparameters. was blended with those of PINN to yield uncertainty estimates for the target outputs assumed to follow a set of PDEs. E-PINN [22] was conjectured to be able to generate uncertainty estimates also for the unknown PDE parameters (apart from the target output) by averaging the PDE residual term over a hypothetical distribution. A limitation of the work of [22] is that it only applies to a subclass of PDEs where its parameters can be separated out as linear terms in a certain form (as expressed in equation 3 of [22]). It was also not clear how one should specify the hypothetical distribution for the learnable PDE parameters.
@@ 2.1 A posterior density from merging principles of Deep Evidential Regression and PINN
where $N_{D}$ is the number of independently sampled points within the domain of the PDE and $\mathcal{R}\left(\partial y,y,x,\vec{\Omega}\right)=0$ refers to the partial differential equation, with $\sigma^{2}_{R}$ being the variance parameter of a density function for the residual. [^2] One can extend this to a system of PDEs by defining a separate residual for each differential equation. The negative logarithm of (2) is the PDE residual loss, with the parameter $\sigma^{2}_{R}$ quantifying the weight of this loss term relative to the data loss term $\sim-\log P(\mathcal{D}|\mathcal{M}(\vec{w}))$. In standard PINN, this is a free parameter and, to our knowledge, there is no principled approach towards determining its choice. Here we lift $\sigma^{2}_{R}$ to be a learnable parameter with the relative weight of the PDE residual evolving as the model shifts towards a minimum in the loss landscape. We regularize the dynamical evolution of $\sigma^{2}_{R}$ through a prior density function $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$ which we pick to be the inverse-gamma distribution. Since the residual is a function of the unknown parameters $\vec{\Omega}$, the form of (1) suggests that we can interpret it as the (unnormalized) posterior density for $\vec{\Omega}$, with $\pi(\vec{\Omega})$ being the prior distribution and the product $P(\mathcal{D}|\mathcal{M}(\vec{w}))\,P(\mathcal{M}(\vec{w})|\vec{\Omega})$ being the likelihood function. The model’s weights are then latent variables with model training equivalent to a maximum a posteriori estimation. This interpretation enables us to compute the uncertainty of $\vec{\Omega}$ as being defined with respect to the posterior density function
@@
where $\gamma$ is the mean output and $\alpha,\beta,\nu$ are learnable functions of the input, related to the overall predictive uncertainty $\sigma^{2}_{p}$ as follows.
@@
where $Y$ denotes each output variable in the dataset $\mathcal{D}$, and the expectation and variance are taken with respect to the t-distribution density function in (6). This distribution is the marginal distribution obtained after integrating the mean and variance of a Gaussian likelihood function $\mathcal{N}(\mu,\sigma^{2})$ with respect to a normal-inverse-gamma prior distribution, where $\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\nu)$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$. We take the dependent variable in the PDE residual to be the mean output, i.e. $y\rightarrow\gamma(x)$ in eqn. (2). This implies that the MAP estimation involves the adherence of the mean target variables to the PDEs. Upon completion of model training, we can place confidence intervals on model’s predictions using eqn. (7). Since we infer $\vec{\Omega}$ at the end of model training via eqn. (3), our framework thus appears as a maximum a posteriori estimation of $\vec{\Omega}$, or more precisely a maximum likelihood estimation that is regularized by the prior $\pi(\vec{\Omega})$.
@@ 2.2 On uncertainty of $\vec{\Omega}$ and its prior distribution
In our framework, we alluded to a posterior density function $f_{p}\left(\vec{\Omega}|\mathcal{D},\mathcal{M}(\vec{w})\right)$ in eqn. (3) of which negative logarithm is the model’s loss function. The uncertainty in $\vec{\Omega}$ reflects the probabilistic deviation of the model from the PDE description, as quantified by the residual. In the following, we will invoke this principle to derive a form for the prior $\pi(\vec{\Omega})$. Let $\mathcal{D}_{\vec{\Omega}}$ denote the finite, discretized domain for the unknown parameters $\vec{\Omega}$. At each point of $\mathcal{D}_{\vec{\Omega}}$, we can evaluate the mean squared deviation between the numerical solution to the PDE characterized by $\vec{\Omega}$ ($L_{p}(x;\vec{\Omega})$) and the data-fitted model $\mathcal{M}_{0}(x;\vec{w}_{0})$.
@@
where $N$ is the normalization constant and all integrals are implemented as numerical Riemann sums over the discretized domain $\mathcal{D}_{\vec{\Omega}}$. We would like the prior distribution of $\Omega$ to be characterized by the same mode and dispersion scales as the highest density region (HDR) of $f$ as discussed in [14]. At some confidence level, say 68%, this region is generally a complex subset of the domain $\mathcal{D}_{\Omega}$. Since our choice of prior distribution affects model training dynamics in the second phase, we adopt a simple Gaussian surrogate distribution for this region, with the means being the modes and the standard deviations being those of each marginal distribution.
@@ 2.3 Determination of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$
we pick its hyperparameters $(\alpha_{r},\beta_{r})$ such that it is consistent with other aspects of our formalism. The mode and mean values are $\frac{\beta_{r}}{\alpha_{r}+1}$ and $\frac{\beta_{r}}{\alpha_{r}-1}$ respectively. Here we restrict ourselves to the case where $\alpha_{r}>1$ so that the mean is well-defined. We pick the initial value of $\sigma^{2}_{R}$ ($\equiv\sigma^{2}_{ini}$) to be the mean of (12). As $\sigma^{2}_{R}$ decreases during model training, it approaches the mode of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$ at which the derivative with respect to $\sigma^{2}_{R}$ vanishes.
@@ 3 Validation case studies: 1D Poisson and 2D Fisher-KPP equations
We validated the E-PINN framework using two representative differential equations : (i)the 1D Poisson equation with Gaussian source (ii)the Fisher-KPP equation. These prototype equations and their variants have found wide applicability across scientific disciplines (see e.g. [6, 10, 13, 19].) For the main performance metric, we adopt the empirical coverage probability (ECP) to measure how well-calibrated the uncertainties are, relative to the nominal ones (NCP). In our case studies, we computed plots of ECP vs NCP and calculated the mean of the absolute discrepancy between these two quantities as a measure of the mean calibration error (MCE) (see e.g. [12, 15]). For assessing if the uncertainty in the inferred parameters $\vec{\Omega}$ is consistent with the model’s deviation from the PDE numerical solutions, we performed a Monte Carlo goodness-of-fit test (see e.g. [24]). This involves estimating an empirical p-value defined as the proportion of posterior samples whose deviation from the posterior mean exceeds that of the model curve.
@@
where $x_{0},\sigma^{2}_{f}$ are parameters to be inferred from the model. The synthetic dataset was based on an exact solution $u_{e}(x)$ defined on the finite segment $x\in[0,1]$, with $x_{0}=1/3,\sigma^{2}_{f}=0.02$. Dirichlet boundary conditions $u=0$ are imposed at $x=0,1$. A uniformly distributed random noise of the form $\sim\mathcal{U}(-0.5,0.5)$ was added to the exact solution within $(0,1)$. We chose $\lambda_{N}$ to be 0.1 of the maximum value of $u_{e}$. The size of the training and testing data was taken to be 240 and 150 (randomly chosen) points respectively covering the entire finite domain. [^3] About half of the training dataset was used as a validation set for choosing the optimal learning rate in the first phase of model training, after which the whole training dataset was used.
For this problem, our base model was a fully-connected two-layer perceptron model with 16 neurons per layer. In the first training phase where the model was fitted purely to data, convergence was attained upon $2\times 10^{5}$ epochs at a learning rate of $10^{-4}$, with the Adam optimizer implemented via PyTorch. Using this initial model, we determined the prior distribution $\pi(\vec{\Omega})$ using eqns. (10) and (11), following our earlier discussion in Section 2.2. Once the prior distribution $\pi(\vec{\Omega})$ was determined, we used it to determine the prior for $\sigma^{2}_{R}$ using eqn.15-18. The finite parameter domains were set to be $x_{0}\in[0,1],\sigma^{2}_{f}\in[0.01,0.06]$ discretized such that the minimal spacing is 1/50 of each interval. We then proceeded to the second phase of model training with a learning rate of $5\times 10^{-4}$. Upon completion of training, our model inferred the equation’s parameters to be
@@
where the upper and lower figures refer to the 68% confidence intervals. We note that they are consistent with the actual values of these parameters ($x_{0}=0.333,\sigma^{2}_{f}=0.02$). The p-value for the Monte Carlo goodness-of-fit test was 0.89, showing lack of evidence against the hypothesis that the model curve arises from the inferred posterior distribution characterized by (19). The model’s uncertainty distribution appeared to be well-calibrated, with the ECP close to the NCP for all confidence intervals. The mean calibration error was 0.02. Figure 1 shows the model-predicted curve and the ECP vs NCP plot revealing a well-calibrated uncertainty distribution.
@@
We set $r=1.6,D=6.2$ for generating the training dataset. These parameters are the ones to be inferred from the model. Like in the Poisson equation case study, we added a uniformly distributed random noise $\sim\mathcal{U}(-0.5,0.5)$ to the exact solution (21) defined on the 2D domain $x\in(-20,20),t\in(0,10)$ (see Fig. 2). The training and testing datasets were set to be $5\times 10^{3}$ and $4\times 10^{3}$ randomly distributed points across the domain. Like in the previous case study, we used a fully connected two-layer perceptron model with 16 neurons per layer. The training dynamics was similar to the previous case, with the model trained for $2\times 10^{5}$ epochs using only the data loss term and the prior distribution constructed from the HDR of the empirical distribution as defined in eqn. (9). The second phase of model training was implemented at a learning rate of $3\times 10^{-4}$, with convergence attained after $3\times 10^{5}$ epochs. At the end of training, our model inferred the equation’s parameters to be
@@
We note that the 68% confidence intervals in (22) are consistent with the actual values $(r=1.6,D=6.2)$. The p-value for the goodness-of-fit test was 0.93, indicating lack of evidence against the posterior distribution of (22). The model’s uncertainty quantification appeared to be well-calibrated similar to the previous Poisson equation case study, with a mean calibration error of 0.024.
@@ 3.3 Comparison against other methods of uncertainty quantification frameworks in PINN
In terms of the degree of uncertainty calibration, we found E-PINN to perform significantly better than Deep Ensemble and B-PINN models for both case studies. While B-PINN tended to yield over-conservative uncertainty estimates, Deep Ensemble models generated excessively over-confident ones. The ECP vs NCP plots showed E-PINN’s uncertainty estimates to be well-calibrated in contrast to the more obvious deviations from the NCP for the other two frameworks. The MCE values for E-PINN were much lower relative to B-PINN and Deep Ensemble. For the 2D Fisher-KPP case study in which heterogeneous noise was inserted into the training data, we found that the E-PINN uncertainty estimates yielded a strong correlation with the training data noise (Spearman’s coefficient $r_{s}\sim 0.8,p<0.01$) which was absent in Deep Ensemble ($\rho_{s}\sim 0.01,p=0.2$) and B-PINN ($\rho_{s}\sim 0.09,p<.001$) as shown in the Fisher-KPP uncertainty plots (Fig. 3).
@@ 4 Application of E-PINN to Bergman minimal model of glucose-insulin regulatory system
where $G$ represents glucose level, $I$ represents insulin concentration and $X$ a more abstract quantity representing glucose uptake activity of insulin-excitable tissue (e.g. liver, adipose, etc.). $G_{b},I_{b}$ are the basal concentrations of glucose and insulin representing baseline values when the effect of the bolus injection of glucose in IVGTT disappears. The parameters $p_{1},p_{2},p_{3}$ represent the glucose decay rate, spontaneous decay rate of remote insulin and the insulin-dependent increase in tissue glucose uptake ability respectively. These are the parameters we leverage E-PINN to infer from measurements of glucose and insulin evolutions in IVGTT samples. In particular, $p_{1}$ is known as an index measuring ‘glucose effectiveness’ (often denoted as $S_{G}$), while $S_{I}\equiv p_{3}/p_{2}$ is known as the insulin sensitivity index that measures the effectiveness at which glucose decreases upon release of insulin. In a IVGTT test, glucose and insulin samples are periodically collected over a period of a few hours. Taking the insulin to be an externally known function, the equations (23), (24) thus constitute a coupled set of first-order ODEs in the variables $G$ and $X$ which is often alluded to as the Bergman minimal model [3].
@@
We apply E-PINN to the modeling of (23) and (24), using the datasets gathered and analyzed in [9, 16, 23]. In particular, we focused on modeling two datasets pertaining to the measured glucose and insulin levels corresponding to healthy and type 2 diabetic subjects. E-PINN enables probabilistic modeling of the datasets and directly yields the posterior distribution of the unknown parameters. Unlike the controlled case studies, we do not have empirical data points for the variable $X(t)$ and thus, our base neural network was designed to yield five outputs – four corresponding to $G(t)$ and its uncertainty, and another corresponding to $X(t)$ so as to avoid overfitting on $X(t)$. We define the PDE residual as the sum of the residual arising from (23) and (24) separately. Estimating $I_{b},G_{b}$ from the datasets, we are left with only $p_{1},p_{2},p_{3}$ as the unknown parameters to be inferred. Fig. 4 depicts our model’s prediction curves and the learned posterior distributions of two clinically useful indices: the glucose effectiveness index $S_{G}$ and insulin sensitivity index $S_{I}$, the latter appearing to be the more discriminatory one, being about 3 orders of magnitude larger for healthy individuals.
@@
Figure 4: The left diagrams show E-PINN’s prediction curves for glucose evolution in healthy and diabetic subjects. Blue dashed curves are the numerical solutions to Bergman’s ODE with parameters $p_{1},p_{2},p_{3}$ being the posterior modes. Deviations between blue and red curves in the middle regions suggest that PDE description may be relatively less valid during those durations. The right diagram superposes the learned posterior distributions of the indices $S_{G},S_{I}$, revealing the latter as a more discriminatory index for identifying diabetic individuals. Log scale for $S_{I}$ was adopted to enable clearer visualization of the distributions for healthy (dark brown) and diabetic (light brown) patients.